OpenStudy (anonymous):

∫ (1/x^2)[cos^2(1/x)]dx Evaluate using substitution method?

OpenStudy (anonymous):

\[\int\limits (1/x^2)[\cos^2(1/x)] dx\]

OpenStudy (anonymous):

put 1/x = t then differentiate this equation to obtain -1/x^2 dx = dt

OpenStudy (anonymous):

annd then? :)

OpenStudy (anonymous):

then replace values in the equation 1/x by t and 1/x^2 by - dt

OpenStudy (anonymous):

express cos^2x in terms of cos x

OpenStudy (anonymous):

then integrate

OpenStudy (anonymous):

cos^2t=(cos2t+1)/2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and after integrating replace t by 1/x

OpenStudy (anonymous):

what happens with the dx? aren't we supposed to replace that too?